### v1

# Injecting Measurement Information Yields a Fast and Noise-Robust Diffusion-Based Inverse Problem Solver

Abstract
Diffusion models have been firmly established as principled zero-shot solvers for linear and nonlinear inverse problems, owing to their powerful image prior and iterative sampling algorithm. These approaches often rely on Tweedie’s formula, which relates the diffusion variate $\mathbf{x}_{t}$ to the posterior mean $\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]$, in order to guide the diffusion trajectory with an estimate of the final denoised sample $\mathbf{x}_{0}$. However, this does not consider information from the measurement $\mathbf{y}$, which must then be integrated downstream. In this work, we propose to estimate the conditional posterior mean $\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t},\mathbf{y}]$, which can be formulated as the solution to a lightweight, single-parameter maximum likelihood estimation problem. The resulting prediction can be integrated into any standard sampler, resulting in a fast and memory-efficient inverse solver. Our optimizer is amenable to a noise-aware likelihood-based stopping criteria that is robust to measurement noise in $\mathbf{y}$. We demonstrate comparable or improved performance against a wide selection of contemporary inverse solvers across multiple datasets and tasks.
1 Introduction
In this work, we study a broad class of problems involving the recovery of a signal $\mathbf{x}$ from a measurement
$\mathbf{y}=\mathcal{A}(\mathbf{x})+\bm{\eta}.$ (1)
with noise $\bm{\eta}$ and measurement operator $\mathcal{A}$. Known as inverse problems, such formulations appear in a multitude of fields, with applications including acoustic reconstruction (Kac 1966), seismic profiling (Hardage 1985), X-ray computed tomography and magnetic resonance imaging (Suetens 2017), and a large number of computer vision reconstruction tasks such as inpainting, deconvolution, colorization, super-resolution, and phase retrieval (Andrews and Hunt 1977).
Figure 1: The posterior mean before and after conditioning on $\mathbf{y}$ for the random inpainting inverse problem.
Figure 2: (a) $\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]$ via the unconditional score versus (b) $\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t},\mathbf{y}]$ via the forward process score estimate obtained by our likelihood maximizer for an image in the FFHQ $256\times 256$ dataset with motion blur applied. With the unconditional score, $\hat{\mathbf{x}}_{0}$ estimates the posterior mean of the dataset, rather than a sample $\mathbf{x}$ that satisfies $\mathcal{A}(\mathbf{x})\approx\mathbf{y}$, especially at $T\gg 0$ (Section 2).
Figure 3: The hazard of over- or under-fitting for a super-resolution task. An ideal noise-aware fit balances between the prior and the noisy measurement $\mathbf{y}$.
Solver Type
No NFE
Backprop
Runtime Memory
DCS (Ours) Hybrid 1x 1x
MCG Projection $2.6\mathbf{x}$ $3.2\mathbf{x}$
DPS Posterior $2.5\mathbf{x}$ $3.2\mathbf{x}$
DPS-JF Posterior $2\mathbf{x} 1.1$$1\mathbf{x}$
DDNM Projection $1.5\mathbf{x}$ $1\mathbf{x}$
[…]
Table 1: Overview of pixel-based solvers used for comparisons in this work. We list the type (Section A.2), whether it requires backpropagation through a neural function evaluation, runtime, and memory footprint.
Often, inverting $\mathcal{A}$ is numerically intractable (Appendix C), meaning that solutions $\mathbf{x}$ satisfying $\mathcal{A}(\mathbf{x})=\mathbf{y}$ are not directly obtainable or unique (Vogel 2002). Moreover, due to measurement noise, it is often possible, but not practically desirable to fit perfectly to $\mathbf{y}$ for risk of overfitting to $\bm{\eta}$ (Aster, Borchers, and Thurber 2018). Therefore, a fundamental step in solving inverse problems is deciding how to select the best option from an equivalence class of solutions, i.e., choosing $\mathbf{x}_{*}\in\{\mathbf{x}:\mathcal{A}(\mathbf{x})\approx\mathbf{y}\}$.
In classical solvers, this is carried out by a regularizer on a normed error loss (Engl, Hanke, and Neubauer 1996):
[…]
where $\epsilon$ is a soft error margin and $R$ is a simple function that satisfies user-specified heuristics, e.g., smoothness or total variation (Beck and Teboulle 2009). However, such approaches often fail to produce realistic results, as it can be difficult to design $R$ without introducing strong biases. With the advent of deep generative models, practitioners found that restricting solutions to the range of a generative model $G$ can greatly improve realism: let $\mathbf{x}=G(\mathbf{w})$ and optimize over $\mathbf{w}$, which can be latent inputs (Bora et al. 2017) or weights (Ulyanov, Vedaldi, and Lempitsky 2018) of a deep neural network. Overall, these methods improve the fidelity of $\mathbf{x}$, but they lack interpretability and require a judiciously selected $R$ and $\epsilon$.
Figure 4: An illustration of our proposed sampling algorithm. An initial noise prediction $\epsilon_{\theta}$ is corrected by the solution $\epsilon_{\mathbf{y}}$ of a noise-aware maximization scheme of the measurement likelihood $p(\mathbf{y}|\mathbf{x}_{t},\epsilon_{y})$. This results in the corrected forward process noise prediction $(\epsilon_{\theta}+\epsilon_{y})\approx-\sigma_{t}^{-1}\nabla\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$. For details see Section 3.
Recently, great strides have been made in solving inverse problems with diffusion models (Ho, Jain, and Abbeel 2020), which produce diverse, realistic samples (Dhariwal and Nichol 2021; Esser et al. 2024) with robust generalization guarantees (Kadkhodaie et al. 2023). Moreover, they are interpretable, directly modeling the (Stein) score $\nabla\log p_{t}$. Unconditional sampling proceeds by reversing a noising process on $\mathbf{x}_{0}\sim p_{\text{data}}$. Solvers then employ a conditional sampling process via a guidance term that pushes samples toward solutions consistent with $\mathbf{y}$:
$\mathbf{x}_{t-1}=\overbrace{\texttt{denoise}[\mathbf{x}_{t},\nabla\log p_{t}(\mathbf{x}_{t})]}^{\texttt{unconditional sampling}}\;+\;\texttt{guidance}$
This approach faces a fundamental challenge: the guidance term depends on a consistency error $\mathcal{A}(\mathbf{x})-\mathbf{y}||$ that is only tractable for $\mathbf{x}=\mathbf{x}_{0}$ (Chung et al. 2022a). Such methods rely (explicitly or implicitly via (Song, Meng, and Ermon 2020)) on Tweedie’s formula, which estimates $\mathbf{x}_{0}$ given a noise prediction $\bm{\epsilon}_{t}\approx-\sigma_{t}\nabla\log p_{t}(\mathbf{x}_{t})$:
$\hat{\mathbf{x}}_{0}=\mathbb{E}\left[\mathbf{x}_{0}|\mathbf{x}_{t}\right]=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\sigma_{t}\bm{\epsilon}_{t}\right).$ (3)
This can then be substituted for $\mathbf{x}$ in the consistency error, producing a differentiable function with respect to $\mathbf{x}_{t}$.
However, a naive implementation of Eq. 3 introduces significant approximation error, as $\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]=\mathbf{x}_{0}$ if and only if $\mathbf{x}_{t}$ is marginally normally distributed (Theorem 2.1). Existing methods use the unconditional score $\nabla\log p_{t}(\mathbf{x}_{t})$ where this assumption does not generally hold. In this work, we instead propose to use the conditional posterior mean $\mathbb{E}\left[\mathbf{x}_{0}|\mathbf{x}_{t},\,\mathbf{y}\right]$ as an estimate for $\mathbf{x}_{0}$ which requires estimating the forward process score $\nabla\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$. While the forward process score is intractable during the reverse process, as $\mathbf{x}_{0}$ is unknown, we utilize the information in $\mathbf{y}$ in a statistically sufficient manner to obtain an estimate (Section 3).
Our contributions are as follows:
  • We identify a limitation with the use of $\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]$ to predict $\mathbf{x}_{0}$ in inverse problems: the approximation is only exact when $\mathbf{x}_{t}$ is normally distributed.
  • We sidestep this by considering the forward process score, $\nabla\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$ which we are able to predict using a maximum likelihood estimator. We show theoretically that our estimate is statistically sufficient to the measurement $\mathbf{y}$, and empirically that it is robust to significant measurement noise (Figure 5).
  • We demonstrate how this score can be plugged into any standard sampler (e.g., DDPM), resulting in an algorithm that is simple, noise-robust, neural backpropagation-free, and stable across time steps. Moreover, it achieves improvements in performance on a large selection of inverse problems, datasets and noise levels.  [^1] Code for method and experiments provided in https://anonymous.4open.science/r/diffusion˙conditional˙sampling.
2 Revisiting Tweedie’s for Inverse Problems
Figure 5: Reconstruction quality at various noise levels $\sigma_{\mathbf{y}}\in\{0.01,0.1,0.5\}$. Our approach strikes a careful balance between quality at each noise level (Table 2) and computational cost (Table 1). More examples in Appendix F.
Diffusion models (Ho, Jain, and Abbeel 2020) reverse a noise-corrupting forward process with marginals $\mathbf{x}_{t}\sim p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$,
$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\underbrace{\sqrt{1-\alpha_{t}}}_{\sigma_{t}}\mathbf{z},\hskip 18.06749pt\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (4)
parameterized by a monotonically time-decreasing scalar $\alpha_{t}$. New samples are generated via the reverse diffusion process which leverages the learned score function $s_{\theta}=-\sigma_{t}^{-1}\bm{\epsilon}_{\theta}\approx\nabla\log p_{t}(\mathbf{x}_{t})$ (Anderson 1982; Vincent 2011; Song et al. 2020). Diffusion-based solvers for inverse problems can be categorized as posterior or projection solvers, and aim to modify the reverse process such that the final variate $\mathbf{x}_{0}$ coincides with a member of the solution set $\{\mathbf{x}_{0}:\mathcal{A}(\mathbf{x}_{0})\approx\mathbf{y}\}$ [^2] We defer a more extensive discussion on diffusion models diffusion-based inverse problem solvers to Appendix A.. This paradigm is afflicted by a fundamental computability paradox: since the consistency error is only explicitly known at $t=0$ via the likelihood function
[…]
we cannot exactly guide the diffusion process at time $t>0$ without first solving for $\mathbf{x}_{0}$. Simultaneously, we cannot generally obtain $\mathbf{x}_{0}$ without first computing $\mathbf{x}_{t}$. Accurately estimating $\hat{\mathbf{x}}_{0}\approx\mathbf{x}_{0}$ is a fundamental problem all solvers must contend with to function properly.
In posterior solvers, the computation of $\nabla\log p(\mathbf{y}|\mathbf{x}_{t})$, is approximated by $\nabla\log p(\mathbf{y}|\hat{\mathbf{x}}_{0})$. In projection solvers, this is hidden in the projection step $\mathbf{P}\mathbf{x}_{t}$, which is driven by a projection on $\hat{\mathbf{x}}_{0}$, followed by a DDIM step (Song, Meng, and Ermon 2020) that involves $\hat{\mathbf{x}}_{0}$. In both cases, Tweedie’s formula is used to create an estimate for $\mathbf{x}_{0}$, given the current $\mathbf{x}_{t}$ by predicting the posterior mean
$\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]=\int_{\mathbb{R}^{d}}\mathbf{x}_{0}p(\mathbf{x}_{0}|\mathbf{x}_{t})d\mathbf{x}_{0}.$ (6)
A surprising limitation of this estimator is that $\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]$ only coincides with $\mathbf{x}_{0}$ when $\mathbf{x}_{t}$ is marginally normally distributed. We formalize this necessary and sufficient condition below:
Theorem 2.1.
Let $\mathbf{x}_{t}$ be sampled from a diffusion process (as in Eq. 4). $\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]=\mathbf{x}_{0}$ if and only if $p(\mathbf{x}_{t})$ is a simple isotropic Gaussian with mean $\sqrt{\alpha_{t}}\mathbf{x}_{0}$ and variance $\sigma_{t}\mathbf{I}$.
Note that $\mathbf{x}_{t}$ is almost never Gaussian, since $\mathbf{x}_{t}$ is distributed as $\phi_{\sigma}\ast p_{\text{data}}$. While $\phi_{\sigma}$ is a simple isotropic Gaussian, $p_{\text{data}}$ is not — the data distribution can be arbitrarily non-convex and multimodal.
We visualize this in Figure 3, where at larger values of $t$, the fidelity of the estimated $\mathbf{x}_{0}$ is poor, resulting in a low quality prediction that is inconsistent with the measurement $\mathbf{y}$. Ultimately, Eq. 6 is a weighted average over all data $\mathbf{x}\sim p_{\text{data}}$, and often cannot properly estimate $\mathbf{x}_{0}$ without incorporating measurement information from $\mathbf{y}$. Instead, we propose to use the conditional posterior mean
$\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t},\mathbf{y}]=\int_{\mathbb{R}^{d}}\mathbf{x}_{0}p(\mathbf{x}_{0}|\mathbf{x}_{t},\mathbf{y})d\mathbf{x}_{0},$ (7)
which only considers those $\mathbf{x}_{0}\sim p_{\text{data}}$ consistent with $\mathbf{y}$. In the following section, we outline a method for directly incorporating this conditional information into the estimate.
Algorithm 1 Diffusion Conditional Sampler (DCS)
1:Input: $\mathbf{y},\mathcal{A},\epsilon_{\theta}$     $|$ Output: $\mathbf{x}_{0}$
[…]
3:for $t=T$ to 11 do
4:  $\bm{\epsilon}\sim\mathcal{N}(\textbf{0},\mathbf{I})$
5:  ${\epsilon_{\mathbf{y}}}\leftarrow\underset{{\epsilon_{\mathbf{y}}}}{\arg\texttt{nam}}\;p_{t}\left(\mathbf{y}|\frac{\mathbf{x}_{t}-\sigma_{t}(\epsilon_{\theta}+\epsilon_{\mathbf{y}})}{\sqrt{\alpha_{t}}}\right)$
6:  $\mathbf{x}_{t-1}\leftarrow\texttt{ddpm\_step}(\mathbf{x}_{t},\epsilon_{\theta}+\epsilon_{y})$
7:end for
Algorithm 2 Noise-aware Maximization (nam)
1:Input: $\mathbf{y},\mathcal{A},\mathbf{x}_{t},\epsilon_{\theta}$$|$ Output: $\epsilon_{y}$
[…]
5:  $\epsilon_{y}\leftarrow\epsilon_{y}+\eta\nabla\log p_{t}\left(\mathbf{y}|\frac{\mathbf{x}_{t}-\sigma_{t}(\epsilon_{\theta}+\epsilon_{\mathbf{y}})}{\sqrt{\alpha_{t}}}\right)$
6:  $\hat{\mathbf{x}}\leftarrow\texttt{Tweedie's}(\mathbf{x}_{t},\epsilon_{\theta}+\epsilon_{y})$
7:end while
3 Diffusion Conditional Sampling
We propose a novel framework for solving inverse problems with diffusion models via a measurement consistent Tweedie’s formula (Eq. 7). At each step, we form a single-parameter measurement model whose maximum likelihood estimator approximates the forward process score $\nabla\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$ (Section 3.1). This estimator is optimized with a noise-robust, likelihood-based early stopping criterion (Section 3.2). The learned score is then input to a standard DDPM sampler (Ho, Jain, and Abbeel 2020), resulting in Algorithm 1. This approach is motivated by both powerful theoretical guarantees (Section 3.3), and yields significant computational advantages (Section 3.4).
3.1 Measurement Likelihood Model
In our setting, we wish to reverse a diffusion process originating from a single fixed $\mathbf{x}_{0}$ — the desired signal $\mathbf{x}$ (Eq. 1), where
$p_{*}(\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\alpha_{t}}\mathbf{x}_{0},(1-\alpha_{t})\mathbf{I}).$ (8)
In standard unconditional sampling, this is also known as the forward process $p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$ (Ho, Jain, and Abbeel 2020). However, when running the reverse process (unconditional sampling) this is a non-trivial quantity to evaluate, as $\mathbf{x}_{0}$ is unknown. We apply Tweedie’s to Eq. 8 and solve for the forward process score via a closed form expression for the likelihood (Eq. 5):
$\log\$$p(\mathbf{y}|\mathbf{x}_{0}({\epsilon_{\mathbf{y}}},\mathbf{x}_{t}))\propto$
[…]
Since $p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$ is distributed as an isotropic Gaussian by construction, knowledge of the true conditional posterior mean would recover $\mathbf{x}_{0}$ exactly (Theorem 2.1). Defining the equivalence,
$\nabla\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})=-\sigma_{t}^{-1}[\epsilon_{\theta}(\mathbf{x}_{t},t)+\epsilon_{\mathbf{y}}],$ (10)
we can solve for our single parameter $\epsilon_{y}$ by maximizing the joint likelihood between the measurement $\mathbf{y}$ and our parameter $\epsilon_{\mathbf{y}}$}. This forms our estimator for $\nabla\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$:
$s_{\text{corrected}}=-\sigma_{t}^{-1}$ $\left[\epsilon_{\theta}(\mathbf{x}_{t},t)+\epsilon_{\mathbf{y}}^{*}\right],$ (11)
where
[…]
Finally, $s_{\text{corrected}}$ can then be input to any standard diffusion model sampler.
Now we turn to estimating $\epsilon_{\mathbf{y}}^{*}$. Given the often noisy and ill-posed nature of Eq. 12 (Appendix C), we seek to select from the solution set $\{\epsilon_{\mathbf{y}}:\mathcal{A}[\hat{\mathbf{x}}_{0}]\approx\mathbf{y}\}$ through a noise-aware maximization algorithm, which we outline below.
Figure 6: DCS performance across several choices of optimizers. LPIPS score of the predicted $\mathbf{x}_{0}$ images is plotted against the natural log of learning rate scaling factor for each optimizer.
3.2 Noise-Aware Maximization
We propose a noise-aware maximization scheme (nam) to improve stability across noise levels. Given a single noisy measurement $\mathbf{y}=\mathcal{A}[\mathbf{x}]+\bm{\eta}$, there is a high risk of overfitting to noise $\bm{\eta}$ (Figures 3 and 5). To mitigate this problem, we propose a maximization scheme with a specialized early stopping criterion based on the measurement likelihood. We leverage the intuition that the corrected forward process score should yield a prediction where a vector of residuals,
$\texttt{res}=\mathbf{y}-\mathcal{A}[\hat{\mathbf{x}}_{0}]$ (13)
should be i.i.d. normally distributed with variance $\sigma_{\mathbf{y}}^{2}$. In other words, each index of res should come from the same distribution as an index of $\bm{\eta}$. Let this be the null hypothesis $\mathbb{H}_{0}$ — we seek to end the likelihood maximization process as soon as $\mathbb{H}_{0}$ holds.
Formally, we optimize Eq. 9 until the likelihood of the alternate hypothesis, $\mathbb{H}_{1}$ (that res is not distributed as $\eta$), is below a desired threshold $p_{\text{critical}}$. Since overfitting is more problematic at the end of sampling ($t\approx 0$) than the beginning of sampling ($t\approx T$), we set $p_{\text{critical}}$ dynamically as a function of $t$, namely $p_{\text{critical}}(t)=\sigma_{t}$. This scheme is heavily inspired by the classical two-sided $\mathbf{z}$-test (Hogg et al. 2013) with $d$ samples, where $d$ is the dimensionality of the image. We use the early-stopping criterion at each time $t$
$\mathbb{P}(|\xi|>|\texttt{res}|\big{|}\mathbb{H}_{0})=2\Phi(-|\texttt{res}|/\sigma_{\mathbf{y}})<\sigma_{t},$ (14)
where $\xi_{i}\overset{\text{iid}}{\sim}\mathcal{N}(\mathbf{0},\sigma_{\mathbf{y}}^{2})$ and $\Phi$ is the CDF of a standard normal distribution. We note that this differs from a classical $\mathbf{z}$-test since we are not seeking to reject the null hypothesis, but optimizing until the null hypothesis can no longer be rejected with sufficiently high probability. The full noise-aware maximization algorithm can be summarized by Algorithm 2. Since our loss function (Eq. 9) is quadratic, our proposed nam has worst-case linear convergence guarantees due to classical results with gradient descent (Boyd and Vandenberghe 2004; Ryu and Boyd 2016) with linear $\mathcal{A}$.
Figure 7: Qualitative comparison of our proposed method against related work on FFHQ $\times256$-1K (left) and ImageNet $\times256$-1K (right). Further comparisons can be found in Appendix F.
3.3 Theory
We highlight two key theoretical properties of our sampler, the correctness of the Tweedie’s approximation and the sufficiency of the resulting score with respect to $\mathbf{y}$.
Correctness
Our algorithm makes use of the conditional variant of Tweedie’s formula, which obeys a very similar set of rules as Tweedie’s (Theorem 2.1).
Theorem 3.1.
Let $\mathbf{x}_{t}$ be sampled from a conditional diffusion process given $\mathbf{y}$ (as in Eq. 4, with $\mathbf{x}_{0}\sim p(\mathbf{x}_{0}|\mathbf{y})$). $\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t},\mathbf{y}]=\mathbf{x}_{0}$ if and only if $p(\mathbf{x}_{t})$ is a simple isotropic Gaussian with mean $\sqrt{\alpha_{t}}\mathbf{x}_{0}$ and variance $\sigma_{t}\mathbf{I}$.
This differs from existing approaches in two key ways. First, we incorporate conditional information from $\mathbf{y}$ directly into the denoising process. Second, rather than $p_{t}(\mathbf{x}_{t})$, we consider $p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$, which is an isotropic Gaussian distribution by construction. Therefore, given $\nabla\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$, Theorem 3.1 tells us that the application of Tweedie’s formula in Eq. 9 will exactly recover $\mathbf{x}_{0}$.
Sufficiency
Even with the optimization framework in Eq. 11, $\nabla\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$ can only be computed up to the information present in $\mathbf{y}$. However, we show that this is provably the best estimate in the following sense: $\epsilon_{\textbf{y}}^{*}$ (and therefore $s_{\text{corrected}}$) is a sufficient statistic for the ground truth $\mathbf{x}_{0}$ given measurement $\mathbf{y}$ under regularity conditions on $\mathcal{A}$ and $\bm{\eta}$:
Theorem 3.2 ($\epsilon_{\mathbf{y}}$$_{*}$ is a sufficient statistic).
Let $\mathbf{y}=\mathcal{A}(\mathbf{x}_{0})+\bm{\eta}$ be an observation from the forward measurement model, and let $\epsilon_{\mathbf{y}}$$_{*}$ be defined as in Eq. 12. Then $p(\mathbf{y}|{\epsilon_{\mathbf{y}}}_{*})=p(\mathbf{y}|\mathbf{x}_{0})$, given that either $\bm{\eta}=0$, or $\mathcal{A}$ is linear.
We extend this result to more general conditions in Theorem B.5. In this sense, DCS effectively closes an information ”leak” by ensuring that the only information about $\mathbf{x}_{0}$ lost in the sampling process is solely that which is irrevocably destroyed by the operator $\mathcal{A}$.
3.4 Computational Efficiency
Empirically, DCS enjoys two main computational advantages. First, it does not need to compute expensive gradients of the score function. Second, it boasts stable performance across choices of $T$ due to the linearity of the forward process diffusion process.
Figure 8: A study on the effect of $T$ on solver performance. DCS method remains nearly invariant to $T$ due to the near-linearity of the forward process diffusion process.
No Expensive $\nabla s_{\theta}(\mathbf{x}_{t},t)$ Evaluations
A drawback of many existing algorithms is the need to compute gradients of the score network during sampling (Table 1). This is the most expensive computation in the diffusion step, increasing the runtime of the algorithm by $\mathbf{2}$-$\mathbf{3}\times$. However, this is unavoidable in these solvers without sacrificing performance (Section 4.2).
A Near-Linear Reverse Process
As DCS models $\mathbb{E}(\mathbf{x}_{0}|\mathbf{x}_{t},\mathbf{y})$, it is able to sample approximately from the reverse diffusion process conditioned on the measurement, which reverses the forward process defined in Eq. 4. If the forward process score was able to be exactly estimated, Tweedie’s recovers $\mathbf{x}_{0}$, and the diffusion process can be solved in a single step. In reality, our approximation of this process is correct up to the information about $\mathbf{x}_{0}$ present in $\mathbf{y}$ (Theorem B.5), under the assumptions detailed in the previous section.
In Figure 8, we validate the robustness of our algorithm to the total diffusion steps ($T$) with the super-resolution task on a subset of the FFHQ $256\times 256$ dataset. We compare against DPS (Chung et al. 2023), DPS-JF (a neural backpropagation-free variant of DPS), and DDNM (Wang, Yu, and Zhang 2022) with additive Gaussian noise at standard deviation $\sigma_{\mathbf{y}}=0.05$.
FFHQ SR $\times 4$ Random Inpainting Box Inpainting Gaussian Deblurring Motion Deblurring
$\sigma_{\mathbf{y}}=0.01$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$
Ours 0.137 30.14 19.45 0.024 34.84 21.19 0.088 25.11 19.25 0.103 28.69 22.62 0.087 29.48 26.67
MCG 0.144 24.84 31.47 0.073 30.59 22.22 0.453 15.44 185.54 0.209 23.51 67.88 0.217 22.93 292.1
DDNM 0.208 26.28 51.33 0.040 33.08 23.35 0.209 18.12 88.32 0.235 26.09 71.47 0.424 14.22 250.9
LGD-MC 0.238 23.45 39.55 0.272 23.46 57.70 0.372 15.45 86.00 0.405 18.78 64.31 0.520 13.90 106.0
DPS 0.163 25.91 33.21 0.105 29.54 29.72 0.113 23.52 24.41 0.129 26.48 26.85 0.159 24.41 29.84
RED-Diff 0.178 29.81 45.68 0.035 33.72 25.03 0.090 25.20 19.98 0.234 29.72 52.09 0.191 29.14 116.9
FFHQ SR $\times 4$ Random Inpainting Box Inpainting Gaussian Deblurring Motion Deblurring
$\sigma_{\mathbf{y}}=0.1$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$
Ours 0.175 24.88 30.11 0.149 27.54 32.80 0.163 23.22 26.44 0.176 25.96 26.08 0.224 24.61 31.40
MCG 0.546 20.44 102.6 0.227 26.00 50.40 0.579 15.30 207.2 0.429 25.80 69.29 0.973 -7.104 295.3
DDNM 0.623 21.49 145.9 0.179 24.96 39.18 0.334 19.20 72.11 1.220 10.73 176.8 0.739 5.099 524.0
LGD-MC 0.256 22.31 39.58 0.288 22.22 56.05 0.384 15.38 87.72 0.415 18.30 66.04 0.524 13.65 105.4
DPS 0.185 24.79 35.46 0.157 26.72 35.24 0.158 22.58 32.47 0.180 24.720 33.53 0.212 22.41 35.09
RED-Diff 0.665 22.10 155.1 0.155 28.62 34.78 0.298 22.96 61.14 0.447 26.93 106.3 0.423 24.16 120.1
ImageNet SR $\times 4$ Random Inpainting Box Inpainting Gaussian Deblurring Motion Deblurring
$\sigma_{\mathbf{y}}=0.01$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$
Ours 0.238 23.45 39.41 0.142 26.06 34.46 0.230 20.63 37.11 0.253 24.22 38.96 0.203 24.62 38.63
MCG 0.638 15.62 89.39 0.198 24.34 35.19 0.273 16.68 80.35 0.645 21.18 124.6 0.980 -5.726 231.1
DDNM 0.333 25.16 51.33 0.084 28.35 20.27 0.258 17.42 85.41 0.456 24.35 67.98 0.694 5.721 304.2
LGD-MC 0.662 14.460 113.5 0.650 14.57 129.8 0.696 11.63 133.9 0.796 10.46 165.7 0.807 9.609 184.7
DPS 0.309 23.99 49.81 0.266 25.05 38.87 0.301 18.76 34.85 0.493 19.14 61.59 0.460 18.65 53.21
RED-Diff 0.386 25.07 57.06 0.090 28.17 16.71 0.239 19.99 54.38 0.459 24.70 68.71 0.376 23.66 55.77
ImageNet SR $\times 4$ Random Inpainting Box Inpainting Gaussian Deblurring Motion Deblurring
$\sigma_{\mathbf{y}}=0.1$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$
Ours 0.402 22.99 48.21 0.166 26.04 34.47 0.243 19.70 46.03 0.407 22.28 51.13 0.435 20.43 61.48
MCG 0.886 14.01 145.1 0.459 19.92 78.86 0.433 15.63 124.0 0.650 22.00 117.4 0.984 -6.868 231.3
DDNM 0.751 20.98 133.3 0.1693 25.63 35.72 0.400 18.06 110.8 1.221 9.602 202.7 0.783 5.009 350.1
LGD-MC 0.671 14.01 116.51 0.661 14.19 131.2 0.701 11.53 134.58 0.804 10.44 167.3 0.806 9.587 185.4
DPS 0.540 18.63 85.06 0.506 20.10 82.74 0.479 18.03 83.06 0.412 20.57 65.07 0.450 18.91 75.65
RED-Diff 0.747 20.66 136.35 0.167 25.38 32.99 0.374 19.68 88.20 0.660 23.19 110.9 0.591 21.27 138.8
Table 2: Quantitative comparison on FFHQ 256x256-1K and ImageNet-1K datasets across various inverse problem tasks and noise levels ($\sigma_{\mathbf{y}}\in\{0.01,0.1\}$).
4 Experiments
We examine the empirical performance of DCS across a variety of natural image based inverse problems, against a range of state-of-the-art methods. Quantitatively, we leverage three key metrics to evaluate the quality of signal recovery: Learned Perceptual Image Patch Similarity (LPIPS), peak signal-to-noise ratio (PSNR), and Frechet Inception Distance (FID).
We run DCS and the other methods listed in Table 1 on the FFHQ-256 (Karras, Laine, and Aila 2019), (Kazemi and Sullivan 2014) and ImageNet (Deng et al. 2009) datasets. For the prior network $\epsilon_{\theta}$, we use the corresponding pretrained model weights from (Chung et al. 2022a).
We examine five operator inversion tasks: Super-Resolution, Gaussian Deblurring, Motion Deblurring, Random Inpainting, and Box Inpainting. We first run experiments with additive Gaussian noise of standard deviation $\sigma_{\mathbf{y}}\in\{0.01,0.1\}$. We also present quantitative results on FFHQ and ImageNet in Table 2, and a qualitative comparison in Figures 7, 5. We delegate further experiments, such as evaluations on subsets of FFHQ used in other works, additional qualitative comparisons, and details of the implementation to Appendix D, E and F.
We find that DCS either outperforms, or is comparable to all existing methods. While some methods may excel at certain metrics in certain tasks and fail to recover the signal at other times, DCS consistently performs well across experiments. For example, DCS is one of few methods that has reasonable results on Motion Deblurring at high noise levels. DDNM and RED-Diff, on the other hand, are powerful across inpainting tasks in general, but fail to perform Motion Deblurring and show underwhelming qualitative performance on many tasks, especially in the high noise regime (Figures 5, 7). We find that projection methods in general overfit in the high noise regime. Other methods such as LGD-MC degrade less severely in the presence of noise, however do not perform as well in lower noise levels indicating underfitting across noise levels. It is apparant from both quantitative and qualitative comparisons that DCS and DPS strike a clearer balance between over- and underfitting.
We also notice that DCS provides a significant speedup and reduction in memory footprint compared to most methods, as noted in Table 1. We achieve this by not requiring backpropagation of the score network, as well as limiting the required number of neural function evaluations by using the measurement-consistent version of Tweedie’s formula.
4.1 Ablation on the Noise-aware Maximization Optimizer
We investigate how the choice of optimizer and parameters affects the noise-aware maximization algorithm in DCS. We note that the flexibility of using an optimizer enables us to make use of the stopping criterion as detailed in Section 3. In Figure 6 we run DCS with AdamW (Loshchilov, Hutter et al. 2017), SGD with momentum, and vanilla SGD to solve the SRx4 task on a subset of FFHQ. Runs of each optimizer at learning rate scaling factors are displayed to show the best performance, ensuring a fair comparison. It is clear in Figure 6 that the addition of a momentum term to the optimization process (both present in AdamW and SGD with momentum) can attain a higher level of image fidelity and solver stability than vanilla SGD. This provides empirical evidence for optimizer bias having an effect on solver performance in DCS. We see from this experiment that AdamW produces the most consistent results across learning rates, which motivates its use in our implementation.
4.2 Improvement on Jacobian-Free Implementations
A major gain in the empirical performance of DCS comes from the fact that it no longer requires backpropagations through the neural score function, which allows for reductions in both runtime and memory footprint. In theory, most diffusion-based solvers can be modified to remove this neural backpropagation step by applying a stop_gradient to the score function output (e.g., RED-Diff (Mardani et al. 2024)). We compare against RED-Diff in the main text, and additionally against ablated variants of DPS (Chung et al. 2022a) and LGD (Song et al. 2023b) in Appendix D.2, and demonstrate clear improvements on these methods.
5 Conclusion
We proposed an effective adjustment to the diffusion-based inverse problem solver framework in the literature that improves speed and stability. Observing that the marginals of the diffusion process which solves the inverse problem is Gaussian distributed at each time $t$, we derived a simple, single-parameter likelihood model, whose sole unknown variate is obtained via a tractable maximum likelihood estimation algorithm. We leveraged this new perspective to create a noise-aware maximization scheme, and demonstrated the effectiveness of our method via a suite of numerical experiments and qualitative comparisons.
Appendix A Background and Related Work
A.1 Diffusion Models
Inspired by non-equilibrium thermodynamics, denoising diffusion probabilistic models (Ho, Jain, and Abbeel 2020) convert data $\mathbf{x}_{0}\sim p_{\text{data}}(\mathbf{x})$ to noise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ via a diffusion process described by the variance-preserving stochastic differential equation (VP-SDE)
$d\mathbf{x}=-\frac{\beta(t)}{2}\mathbf{x}dt+\sqrt{\beta(t)}d{\mathbf{w}},$ (15)
where $\beta(t):\mathbb{R}\rightarrow[0,1]$ is a monotonically increasing noise schedule and $\mathbf{w}$ is the standard Wiener process (Song et al. 2020). This leads to the marginal distribution
$p_{t}(\mathbf{x}_{t})=\mathbb{E}_{\mathbf{x}_{0}\sim p_{\text{data}}}\big{[}\mathcal{N}(\mathbf{x}_{t};\sqrt{\alpha_{t}}\mathbf{x}_{0},\;\underbrace{(1-\alpha_{t})}_{\sigma_{t}^{2}}\mathbf{I})\big{]},\hskip 7.22743pt\alpha_{t}=e^{-\frac{1}{2}\int_{0}^{t}\beta(s)ds},$ (16)
where $\mathcal{N}(\ \cdot\ ;\mu,\Sigma)$ is the probability density function (pdf) of a normal distribution centered at $\mu$ with covariance $\Sigma$. Sampling from $p_{\text{data}}(\mathbf{x})$ can then occur by modeling the reverse diffusion, which has a simple form given by (Anderson 1982)
$d\overline{\mathbf{x}}=\left[-\frac{\beta(t)}{2}\mathbf{x}-\beta(t)\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x}_{t})\right]dt+\sqrt{\beta(t)}d\overline{\mathbf{w}},$ (17)
with reverse-time Wiener process $\overline{\mathbf{w}}$ and score function $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x}_{t})$. Therefore, diffusion model training consists of approximating the score function with a model
$s_{\theta}(\mathbf{x}_{t},t)\approx\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x}_{t}),$ (18)
and sampling consists of obtaining solutions to the reverse-time SDE (17) with numerical solvers. A simple approach is given by the DDIM sampler with $\sigma_{t}=\sqrt{1-\alpha_{t}}$ (Song, Meng, and Ermon 2020)
$\mathbf{x}_{t-1}=\sqrt{\alpha_{t-1}}\frac{\mathbf{x}_{t}+\sigma^{2}_{t}\nabla\log p_{t}(\mathbf{x}_{t})}{\sqrt{\alpha_{t}}}+\sigma_{t-1}\bm{\epsilon},$ (19)
where $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$.
A.2 Solving Inverse Problems with Diffusion Models
When solving inverse problems with diffusion models, the aim is to leverage information from $\mathbf{y}$ to define a modified reverse diffusion process
$\mathbf{x}_{T},\mathbf{x}_{T-1},\dots,\mathbf{x}_{1},\mathbf{x}_{0},$ (20)
such that $\mathbf{x}_{t}$ coincides with the desired $\mathbf{x}$ (Eq. 1) precisely at $t=0$. Previous approaches can generally be sorted into two categories, which we designate posterior solvers and projection solvers.
Posterior Solvers
An intuitive approach is leveraging Bayes’ rule to sample from the posterior distribution given a prior $p_{t}(\mathbf{x}_{t})$ and observation $\mathbf{y}$:
$\mathbf{x}_{t}\sim p(\mathbf{x}_{t}|\mathbf{y})=\frac{p(\mathbf{y}|\mathbf{x}_{t})p(\mathbf{x}_{t})}{p(\mathbf{y})}.$ (21)
Taking logs and gradients of both sides of the equation, we obtain a form of the conditional density that can be accurately approximated with the modeled score function
$\nabla\log p(\mathbf{x}_{t}|\mathbf{y})=\nabla\log p(\mathbf{y}|\mathbf{x}_{t})+\nabla\log p(\mathbf{x}_{t})\approx\nabla\log p(\mathbf{y}|\mathbf{x}_{t})+s_{\theta}(\mathbf{x}_{t},t),$ (22)
and describes the core method of the DPS algorithm (Chung et al. 2022a). This strategy can also be extended to latent diffusion models, resulting in Latent-DPS and PSLD (Rout et al. 2023). Generally, the conditional term $\nabla\log p(\mathbf{y}|\mathbf{x}_{t})$ cannot be exact due to reasons we will investigate subsequently in Section 2, though these approximations are improved in LGD (Song et al. 2023b) and STSL (Rout et al. 2024). More recent work (Sun et al. 2024) propose an annealed Monte-Carlo-based perspective to posterior sampling, which results in a very similar algorithm to DPS. Much like MCG and ReSample (discussed in the next category), posterior solvers require estimating $\frac{\partial}{\partial\mathbf{x}_{t}}\mathbf{x}_{0}$ which involves backpropagation through the diffusion model, and significantly increases runtime and hampers scalability compared to unconditional sampling.
Projection Solvers
Another approach involves guiding the reverse diffusion process by directly projecting $\mathbf{x}_{t}$ onto a manifold $\mathcal{M}=\{\mathbf{x}:\mathcal{A}(\mathbf{x})=y\}\subseteq\mathbf{R}^{d}$ at each time step, i.e.
$\mathbf{x}_{t}^{\prime}$ $=\mathbf{P}\hat{\mathbf{x}}_{0}[\mathbf{x}_{t}]$ (23)
$\mathbf{x}_{t-1}$ $=\sqrt{\alpha_{t-1}}\frac{\mathbf{x}_{t}^{\prime}+\sigma^{2}_{t}\nabla\log p(\mathbf{x}_{t}^{\prime}|\hat{\mathbf{x}}_{0}[\mathbf{x}_{t}])}{\sqrt{\alpha_{t}}}+\sigma_{t-1}\bm{\epsilon}.$ (24)
Where $\hat{\mathbf{x}}_{0}[\mathbf{x}_{t}]$ is some prediction of $\mathbf{x}_{0}$ given only $\mathbf{x}_{t}$ (we elaborate in Section 2), and $\mathbf{P}$ is either a projection onto the low rank subspace or range of $\mathcal{A}$. The resulting algorithms are DDRM (Kawar et al. 2022) and DDNM (Wang, Yu, and Zhang 2022), respectively. Of course, this strategy is often restricted to situations where two conditions simultaneously hold true: (1) the measurement operator $\mathcal{A}$ is linear, and (2) the inverse problem is noiseless, i.e, $\eta$ is identically $\mathbf{0}$. These assumptions drastically limit the applicability of such models. The linearity restriction can be lifted by taking derivatives the measurement discrepancy, as in MCG (Chung et al. 2022b) and ReSample (Song et al. 2024), though this comes at the cost of significantly increased computation, requiring $\frac{\partial}{\partial\mathbf{x}_{t}}\mathbf{x}_{0}$ which involves backpropagating through the score network. Finally, (Cardoso et al. 2023) straddles the line between both categories — while MCGdiff is ostensibly a Bayesian solver, it bears greater resemblance to projection solvers since it does not form the decomposition in Eq. 22 and also samples by projecting each iterate to the null-space of $\mathcal{A}$, thus implementing a projected n-particle sequential monte carlo (SMC) sampling algorithm.
A Maximum Likelihood Solver
We take a different perspective on solving the inverse problem. As seen in Section 2, both projection and posterior solvers must quantify the discrepancy between $\mathbf{x}_{t}$ and $\mathbf{y}$ via the consistency error $\mathcal{A}(\mathbf{x}_{0})-\mathbf{y}||$ at each diffusion step. Due to the complexity of the diffusion process, this involves approximating a fundamentally intractable quantity. In Section 3, we construct a simpler process whose parameters can be obtained via maximum likelihood estimation. Unlike the evidence lower bound proposed in (Mardani et al. 2024), we derive an explicit likelihood model, which is amenable to an optimization scheme with a probabilistic noise-aware stopping criterion. Finally, we show that the resulting algorithm is simple, fast, and adaptable to noise.
Appendix B Additional Theorems and Proofs
B.1 Proof of Tweedie’s Formula
For completeness, we include the statement and proof for Tweedie’s formula.
Theorem B.1 (Tweedie’s Formula).
Let $\mathbf{x}_{0}$ be a sample drawn from a distribution $p(\mathbf{x}_{0})$. Then for any
$\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\mathbf{z}\hskip 36.135pt\mathbf{z}\sim\mathcal{N}(0,\mathbf{I})$ (25)
drawn from the marginal of the diffusion process on $p(\mathbf{x}_{0})$ at time $t$, the posterior mean given $\mathbf{x}_{t}$ is
$\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]=\frac{1}{\alpha_{t}}\left[\mathbf{x}_{t}+\sigma_{t}^{2}\nabla\log p_{t}(\mathbf{x}_{t})\right].$ (26)
Proof (of Lemma 26).
Let $\phi_{\sigma}$ be the pdf of $\mathcal{N}(0,\sigma\mathbf{I})$. We first note that the marginal distribution at time $t$ can be written as
$p_{t}(\mathbf{x}_{t})=\left(p_{\alpha_{t}}\ast\phi_{\sigma_{t}}\right)(\mathbf{x}_{t})=\int\phi_{\sigma_{t}}(\mathbf{x}_{t}-\mathbf{x})p_{\alpha_{t}}(\mathbf{x})d\mathbf{x},$ (27)
where
$p_{\alpha_{t}}(\mathbf{x}|\mathbf{y})=\frac{1}{\alpha_{t}}p\left(\alpha_{t}^{-1}\mathbf{x}\right)$ (28)
due to the probabilistic change-of-variables formula. Letting $\mathbf{x}^{\prime}=\alpha_{t}\mathbf{x}_{0}$, we have the equality
$\frac{\mathbb{E}[\mathbf{x}^{\prime}|\mathbf{x}_{t},\mathbf{y}]-\mathbf{x}_{t}}{\sigma_{t}^{2}}$ $=\int\frac{\mathbf{x}^{\prime}-\mathbf{x}_{t}}{\sigma_{t}^{2}}p(\mathbf{x}^{\prime}|\mathbf{x}_{t})d\mathbf{x}^{\prime}$
$=\int\frac{\mathbf{x}^{\prime}-\mathbf{x}_{t}}{\sigma_{t}^{2}}\frac{p(\mathbf{x}^{\prime},\mathbf{x}_{t})}{p(\mathbf{x}_{t})}d\mathbf{x}^{\prime}$
$=\int\frac{\alpha_{t}\mathbf{x}_{0}-\mathbf{x}_{t}}{\sigma_{t}^{2}}\frac{\phi_{\sigma_{t}}(\mathbf{x}_{t}-\mathbf{x}^{\prime})p_{\alpha_{t}}(\mathbf{x}^{\prime})}{\int\phi_{\sigma_{t}}(\mathbf{x}_{t}-\mathbf{x})p_{\alpha_{t}}(\mathbf{x})d\mathbf{x}}d\mathbf{x}^{\prime}$
$=\int\left[\nabla_{\mathbf{x}_{t}}\phi_{\sigma_{t}}(\mathbf{x}_{t}-\mathbf{x}^{\prime})\right]\frac{\phi_{\sigma_{t}}(\mathbf{x}_{t}-\mathbf{x}^{\prime})p_{\alpha_{t}}(\mathbf{x}^{\prime})}{\int\phi_{\sigma_{t}}(\mathbf{x}_{t}-\mathbf{x})p_{\alpha_{t}}(\mathbf{x})d\mathbf{x}}d\mathbf{x}^{\prime}$
$=\nabla_{\mathbf{x}_{t}}\log\left[\phi_{\sigma_{t}}(\mathbf{x}_{t}-\mathbf{x}^{\prime})p_{\alpha_{t}}(\mathbf{x}^{\prime})\right]$
$=\nabla\log p_{t}(\mathbf{x}_{t}).$
Re-arranging terms on either side of the equation, we obtain
$\mathbb{E}[\mathbf{x}^{\prime}|\mathbf{x}_{t}]=\mathbf{x}_{t}+\sigma_{t}^{2}\nabla\log p_{t}(\mathbf{x}_{t}).$ (29)
Finally, we expand $\mathbf{x}^{\prime}=\alpha_{t}\mathbf{x}_{0}$ and invoke the linearity of the expectation to arrive at
$\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]=\frac{1}{\alpha_{t}}\left[\mathbf{x}_{t}+\sigma_{t}^{2}\nabla\log p_{t}(\mathbf{x}_{t}|\mathbf{y})\right].$ (30)
as desired. ∎
B.2 Proof for Theorem 2.1 ($\mathbf{x}_{t}\sim\mathcal{N}(\bm{\mu}(\mathbf{x}_{0}),\sigma_{t}^{2}\mathbf{I})\iff\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]=\mathbf{x}_{0}$)
We demonstrate sufficiency of the Gaussian-distributed condition by proving Lemma B.2.
Lemma B.2 (Sufficent condition).
Let $\mathbf{x}_{0}$ be given. Suppose $\mathbf{x}_{t}$ is distributed as
$p_{t}(\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\alpha_{t}}\mathbf{x}_{0},\underbrace{1-\alpha_{t}}_{\sigma_{t}^{2}}\mathbf{I}).$ (31)
Then $\mathbf{x}_{0}$ can be recovered via
$\mathbf{x}_{0}=\frac{1}{\sqrt{\alpha_{t}}}\left[\mathbf{x}_{t}+\sigma_{t}^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})\right].$ (32)
Proof (of Lemma B.2).
[…]
$=\frac{1}{\sqrt{\alpha_{t}}}\left[\mathbf{x}_{t}-(\mathbf{x}_{t}-\sqrt{\alpha_{t}}\mathbf{x}_{0})\right]$ (34)
$=\mathbf{x}_{0}.$ (35)
To demonstrate the necessary condition, we show that the inverse of Lemma B.2 also holds.
Lemma B.3 (Necessary condition).
If $\mathbf{x}_{0}$ can be recovered via Eq. 32, then $p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$ takes the form Eq. 31.
Proof (of Lemma B.3).
Suppose that
$\mathbf{x}_{0}=\frac{1}{\sqrt{\alpha_{t}}}\left[\mathbf{x}_{t}+\sigma_{t}^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})\right]$ (36)
Then we may re-arrange terms, obtaining
$\frac{\sqrt{\alpha_{t}}\mathbf{x}_{0}-\mathbf{x}_{t}}{\sigma_{t}^{2}}=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}).$ (37)
Taking the anti-derivative of both sides, we conclude that
[…]
Since $\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$ can only take this form when $p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$ is a simple isotropic Gaussian distribution, we conclude our proof. ∎
Proof (of Theorem 2.1).
First, we note that Tweedie’s formula (Efron 2011) tells us that the posterior mean of a data distribution $\mathbf{x}_{t}\sim p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$ can be obtained via the relation
$\mathbf{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]=\frac{1}{\sqrt{\alpha_{t}}}\left[\mathbf{x}_{t}+\sigma_{t}^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})\right].$ (39)
Then, since Lemmas B.2 and B.3 are converses of each other, we demonstrate that the conditions stated in Lemma B.2 are necessary and sufficient. ∎
B.3 Proof for Theorem 3.1 ($\mathbf{x}_{t}\sim\mathcal{N}(\bm{\mu}(\mathbf{x}_{0},\mathbf{y}),\sigma_{t}^{2}\mathbf{I})\iff\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t},\mathbf{y}]=\mathbf{x}_{0}$)
First, we show that the measurement consistent Tweedie’s formula holds for a given diffusion variate $\mathbf{x}_{t}$ and measurement $\mathbf{y}$ .
Theorem B.4 (Conditional Tweedie’s Formula).
Let $\mathbf{x}_{0}$ be a sample drawn from a conditional distribution $p(\mathbf{x}_{0}|\mathbf{y})$. Then for any
$\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\mathbf{z}\hskip 36.135pt\mathbf{z}\sim\mathcal{N}(0,\mathbf{I})$ (40)
drawn from the marginal of the diffusion process on $p(\mathbf{x}_{0}|\mathbf{y})$ at time $t$, the conditional posterior mean given $\mathbf{x}_{t}$ is
$\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t},\mathbf{y}]=\frac{1}{\alpha_{t}}\left[\mathbf{x}_{t}+\sigma_{t}^{2}\nabla\log p_{t}(\mathbf{x}_{t}|\mathbf{y})\right].$ (41)
Proof (of Lemma 41).
Let $\phi_{\sigma}$ be the pdf of $\mathcal{N}(0,\sigma\mathbf{I})$. We first note that the marginal distribution at time $t$ can be written as
$p_{t}(\mathbf{x}_{t}|\mathbf{y})=\left(p_{\alpha_{t}}(\cdot|\mathbf{y})\ast\phi_{\sigma_{t}}\right)(\mathbf{x}_{t})=\int\phi_{\sigma_{t}}(\mathbf{x}_{t}-\mathbf{x})p_{\alpha_{t}}(\mathbf{x}|\mathbf{y})d\mathbf{x},$ (42)
where
$p_{\alpha_{t}}(\mathbf{x}|\mathbf{y})=\frac{1}{\alpha_{t}}p\left(\alpha_{t}^{-1}\mathbf{x}|\mathbf{y}\right)$ (43)
due to the probabilistic change-of-variables formula. Letting $\mathbf{x}^{\prime}=\alpha_{t}\mathbf{x}_{0}$, we have the equality
$\frac{\mathbb{E}[\mathbf{x}^{\prime}|\mathbf{x}_{t},\mathbf{y}]-\mathbf{x}_{t}}{\sigma_{t}^{2}}$ $=\int\frac{\mathbf{x}^{\prime}-\mathbf{x}_{t}}{\sigma_{t}^{2}}p(\mathbf{x}^{\prime}|\mathbf{x}_{t},\mathbf{y})d\mathbf{x}^{\prime}$
$=\int\frac{\mathbf{x}^{\prime}-\mathbf{x}_{t}}{\sigma_{t}^{2}}\frac{p(\mathbf{x}^{\prime},\mathbf{x}_{t}|\mathbf{y})}{p(\mathbf{x}_{t}|\mathbf{y})}d\mathbf{x}^{\prime}$
$=\int\frac{\alpha_{t}\mathbf{x}_{0}-\mathbf{x}_{t}}{\sigma_{t}^{2}}\frac{\phi_{\sigma_{t}}(\mathbf{x}_{t}-\mathbf{x}^{\prime})p_{\alpha_{t}}(\mathbf{x}^{\prime}|\mathbf{y})}{\int\phi_{\sigma_{t}}(\mathbf{x}_{t}-\mathbf{x})p_{\alpha_{t}}(\mathbf{x}|\mathbf{y})d\mathbf{x}}d\mathbf{x}^{\prime}$
$=\int\left[\nabla_{\mathbf{x}_{t}}\phi_{\sigma_{t}}(\mathbf{x}_{t}-\mathbf{x}^{\prime})\right]\frac{\phi_{\sigma_{t}}(\mathbf{x}_{t}-\mathbf{x}^{\prime})p_{\alpha_{t}}(\mathbf{x}^{\prime}|\mathbf{y})}{\int\phi_{\sigma_{t}}(\mathbf{x}_{t}-\mathbf{x})p_{\alpha_{t}}(\mathbf{x}|\mathbf{y})d\mathbf{x}}d\mathbf{x}^{\prime}$
$=\nabla_{\mathbf{x}_{t}}\log\left[\phi_{\sigma_{t}}(\mathbf{x}_{t}-\mathbf{x}^{\prime})p_{\alpha_{t}}(\mathbf{x}^{\prime}|\mathbf{y})\right]$
$=\nabla\log p_{t}(\mathbf{x}_{t}|\mathbf{y}).$
Re-arranging terms on either side of the equation, we obtain
$\mathbb{E}[\mathbf{x}^{\prime}|\mathbf{x}_{t},\mathbf{y}]=\mathbf{x}_{t}+\sigma_{t}^{2}\nabla\log p_{t}(\mathbf{x}_{t}|\mathbf{y}).$ (44)
Finally, we expand $\mathbf{x}^{\prime}=\alpha_{t}\mathbf{x}_{0}$ and invoke the linearity of the expectation to arrive at
$\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t},\mathbf{y}]=\frac{1}{\alpha_{t}}\left[\mathbf{x}_{t}+\sigma_{t}^{2}\nabla\log p_{t}(\mathbf{x}_{t}|\mathbf{y})\right].$ (45)
as desired. ∎
Now, the main theorem follows.
Proof (of Theorem 2.1).
We note that the proofs for Lemmas B.2 and B.3 remain the same if we let $p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})=p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{y})$. Thus we again observe that the proofs for Lemmas B.2 and B.3 are converses of each other, and demonstrate that the conditions stated in Lemma B.2 are necessary and sufficient. ∎
B.4 Theorems for Sufficiency
We set up Theorems to show that the estimator in Eq. 11 is a sufficient statistic under different properties of $\mathcal{A}$. Letting $\mathbf{f}(\mathbf{y})$ be the function that obtains $\nabla\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$ via Eq. 11, we show that $\mathbf{y}$ is measurable under the sigma algebra induced by the measurement $\mathbf{f}$.
Intuitively, we demonstrate that $\mathbf{f}(\mathbf{y})$ contains as much information as possible about the underlying signal $\mathbf{x}_{0}$ as can be gathered via $\mathbf{y}$. The theoretical and intuitive statements can be summarized by the simple conditional equivalence
$p(\mathbf{y}|{\epsilon_{\mathbf{y}}}_{*})=p(\mathbf{y}|\mathbf{x}_{0}).$ (46)
In Theorem 3.2, we consider two simple and theoretically similar cases: when $\mathbf{y}=\mathcal{A}(\mathbf{x})$ is noise-free, and when $\mathcal{A}$ is linear. We restate it here in a less condensed form for clarity:
Theorem 3.2. Let $\mathbf{y}=\mathcal{A}(\mathbf{x}_{0})+\bm{\eta}$ be an observation from the forward measurement model, and let
${\epsilon_{\mathbf{y}}}_{*}=\underset{\epsilon_{\mathbf{y}}}{\text{argmax}}\quad\log p\left(\mathbf{y}\bigg{|}\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{x}_{t}+\sigma^{2}_{t}{\epsilon_{\mathbf{y}}})\right).$ (47)
Then
$p(\mathbf{y}|{\epsilon_{\mathbf{y}}}_{*})=p(\mathbf{y}|\mathbf{x}_{0}).$ (48)
given that either $\bm{\eta}=0$, or $\mathcal{A}$ is linear.
While we do not evaluate the empirical performance of DCS on non-linear operators, we additionally investigate the general noisy case where $\mathcal{A}$ is allowed to be nonlinear. We find that our results can still be quite general: we only need to assume $\mathcal{A}$ surjective, meaning that there exists some $\mathbf{x}\in\text{domain}(\mathcal{A})$ such that $\mathcal{A}(\mathbf{x})=\mathbf{y}$. In fact, this result is slightly stronger — we are able to show that sufficiency holds for $\mathcal{A}$ that are compositions of linear and surjective functions.
Theorem B.5.
Let $\epsilon_{\mathbf{y}}$$_{*}$ be as defined in Theorem 3.2. Suppose the twice-differentiable operator $\mathcal{A}:=\mathbf{P}^{T}\circ\mathbf{\phi}$ is composed of $\mathbf{P}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{r}$, a linear projection, and $\mathbf{\phi}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{r}$, an arbitrary surjective function. We have that
$p(\mathbf{y}|{\epsilon_{\mathbf{y}}}_{*})=p(\mathbf{y}|\mathbf{x}_{0}).$ (49)
To prove Theorems 3.2 and B.5, we establish the following Lemma which characterizes useful information about $\mathbf{x}_{0}^{*}$.
Lemma B.6.
Suppose $\mathbf{y}\in\mathbb{R}^{k}$ is fixed, $\mathbf{x}_{t}\in\mathbb{R}^{n}$, with twice differentiable linear operator $\mathcal{A}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{k}$. Then, for ${\epsilon_{\mathbf{y}}}=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$ which maximizes $p(\mathbf{y}|\mathbf{x}_{0})$, the following holds true:
  1. 1.
    if $\eta=0$ (i.e. the noiseless regime), $\mathcal{A}(\mathbf{x}_{0})=\mathcal{A}(\mathbf{x}_{t}+\sigma_{t}^{2}{\epsilon_{\mathbf{y}}}^{*})$
  2. 2.
    if $\mathcal{A}$ is surjective, $\mathcal{A}(\mathbf{x}_{0})=\mathcal{A}(\mathbf{x}_{t}+\sigma_{t}^{2}{\epsilon_{\mathbf{y}}}^{*})$
  3. 3.
    if $\mathcal{A}$ is linear, $\langle\mathbf{y}-\mathcal{A}(\mathbf{x}_{t}+\sigma_{t}^{2}{\epsilon_{\mathbf{y}}}^{*}),\mathcal{A}(\mathbf{x}_{t}+\sigma_{t}^{2}{\epsilon_{\mathbf{y}}}^{*})-\mathcal{A}(\mathbf{x}_{0})\rangle=0$.
An interpretation of statement 3 reads that the optimal solution $\epsilon_{\mathbf{y}}$$^{*}$ for estimating $\mathbf{x}_{0}$ is orthogonal to the error to $\mathbf{y}$ in the linear case. The requirements of statement 3 may be relaxed to the statement $\mathcal{A}(\mathbf{x})-\mathcal{A}(\mathbf{z})$ is in the range of the Jacobian of $\mathcal{A}$ at $\mathbf{z}$, however this is less intuitive than linearity. We avoid invoking linearity of $\mathcal{A}$ as long as possible to illustrate the fact that other transformations may share this property as well.
Proof (of Lemma B.6).
We will make use of the bijective mapping $\mathbf{z}\mapsto\mathbf{x}_{t}+\sigma_{t}^{2}{\epsilon_{\mathbf{y}}}$, and charactarize the minima which maximize $\log p(\mathbf{y}|\mathbf{x}_{0})$. We can solve the optimization problem,
[…]
A minima to this objective can be characterized by the first order necessary condition,
[…]
$=-2\mathbf{J}_{\mathbf{z}}[\mathcal{A}](\mathbf{z})^{T}(\mathcal{A}(\mathbf{x}_{0})-\eta-\mathcal{A}(\mathbf{z})):=0.$
We can confirm it is a minima by checking the solution of the above with,
[…]
$=2\mathbf{J}_{\mathbf{z}}[\mathcal{A}](\mathbf{z}^{*})^{T}\mathbf{J}_{\mathbf{z}}[\mathcal{A}](\mathbf{z}^{*})+\sum_{j=1}^{k}\mathbf{H}_{\mathbf{z}}[\mathcal{A}_{(j)}](\mathbf{z}^{*})\left(\mathbf{y}-\mathcal{A}(\mathbf{z}^{*})\right)$
$\succcurlyeq 0.$
If $\eta=0$, we have that $\mathcal{A}(\mathbf{x}_{0})=\mathbf{y}$, and therefore choosing any $\mathcal{A}(\mathbf{z}^{*})=\mathcal{A}(\mathbf{x}_{0})$ satisfies the first order condition. The second order condition is furthermore satisfied, as $\mathbf{y}-\mathcal{A}(\mathbf{z}^{*})=0$, meaning,
[…]
This satisfies statement 1. Statement 2 is satisfied similarly, by choosing the same $\mathbf{z}$. Note that this case differs, in that $\mathbf{z}=\mathbf{x}_{0}$ is no longer necessarily a valid solution.
Statement 3, is already satisfied in the cases where $\mathcal{A}$ has rank equal to the dimension of its co-domain (if $d=n$, this is equivalent to being full rank), since $\mathbf{y}-\mathcal{A}(\mathbf{z}^{*})=0$. Therefore we assume $\mathcal{A}$ is low-rank to prove the remaining cases.
[…]
$\delta_{\perp}^{*}$ $=\mathbf{y}-\mathcal{A}(\mathbf{z}^{*})$
$=\mathcal{A}(\mathbf{x}_{0})+\eta-\mathcal{A}(\mathbf{z}^{*})$
$=\mathcal{A}(\mathbf{x}_{0})+\delta_{\perp}^{*}+\delta^{*}-\mathcal{A}(\mathbf{z}^{*})$
$\implies\delta^{*}$ $=\mathcal{A}(\mathbf{z}^{*})-\mathcal{A}(\mathbf{x}_{0}).$
At the optima of the original objective, $\mathbf{z}^{*}$, the first order necessary condition dictates that,
$\mathbf{J}_{\mathbf{z}}[\mathcal{A}](\mathbf{z}^{*})^{T}(y-\mathcal{A}(\mathbf{z}^{*}))$ $=\mathbf{J}_{\mathbf{z}}[\mathcal{A}](\mathbf{z}^{*})^{T}\delta_{\perp}^{*}:=0.$
For a linear $\mathcal{A}$, the Jacobian is constant, so let $\mathbf{J}_{\mathbf{z}}[\mathcal{A}]=\mathbf{J}$. Therefore, $\mathbf{J}^{T}\delta_{\perp}^{*}=0$, meaning $\delta_{\perp}^{*}\in\mathcal{N}(\mathbf{J}^{T})$.
Simultaneously, since $\delta^{*}=\mathcal{A}(\mathbf{x})-\mathcal{A}(\mathbf{z}^{*})=\mathcal{A}(\mathbf{x}-\mathbf{z}^{*})=\mathbf{J}(\mathbf{x}-\mathbf{z}^{*})$, we have $\delta^{*}\in\mathcal{R}(\mathbf{J}^{T})$. Therefore due to the orthogonality of range and null spaces of a matrix, $\langle\delta_{\perp}^{*},\,\delta^{*}\rangle=0$, completing the proof.
We are now able to prove the theorems in the main text.
Proof of Theorem 3.2.
We leverage the theory of sufficient statistics to demonstrate our result. Namely, if $\epsilon_{\mathbf{y}}$$_{*}$ is a sufficient statistic for $\mathbf{y}$, then,
$p(\mathbf{y}|{\epsilon_{\mathbf{y}}}_{*})=p(\mathbf{y}|{\epsilon_{\mathbf{y}}}_{*},\mathbf{x}_{0})=p(\mathbf{y}|\mathbf{x}_{0}).$ (50)
Therefore it suffices to demonstrate that $\epsilon_{\mathbf{y}}$$_{*}$ is a sufficient statistic for $\mathbf{y}$.
By the Neyman-Fisher Factorization theorem, we have that a necessary and sufficient condition is if there exists non-negative functions $g_{\theta}$ and $h$ such that
$p(\mathbf{y}|\mathbf{x}_{0})=g({\epsilon_{\mathbf{y}}}_{*},\mathbf{x}_{0})h(\mathbf{y}).$ (51)
We observe that since $\eta\sim\mathcal{N}(\mathbf{0},\sigma_{\mathbf{y}}^{2}\mathbf{I})$, our random variable $\mathbf{y}$ can be characterized by the density function
$p(\mathbf{y}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{y};\mu=\mathcal{A}(\mathbf{x}_{0}),\Sigma=\sigma_{\mathbf{y}}^{2}\mathbf{I}).$ (52)
Therefore, letting $\mathbf{y}_{{\epsilon_{\mathbf{y}}}_{*}}=\mathcal{A}(\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{x}_{t}+\sigma^{2}_{t}{\epsilon_{\mathbf{y}}}_{*}))$, we can write
[…]
where the third equality is due to Lemma B.6. In the case that $\mathcal{A}$ is surjective, or the noiseless regime, statements 2 and 1 respectively satisfy the equality above trivially, as $\mathbf{y}=\mathbf{y}_{{\epsilon_{\mathbf{y}}}_{*}}$. If the operator is otherwise linear, statement 3 shows the cross term vanishes.
Therefore, we can assign
[…]
In the case where the measurement process $\mathcal{A}(\mathbf{x})=\mathbf{y}$ is noiseless, this implies $h(\mathbf{y})=1$. ∎
We now modify the argument in order to relax the linearity assumption.
Proof of Theorem B.5.
[…]
Since $\mathbf{z}^{*}$ minimizes the objective $\mathbf{y}-\mathcal{A}(\mathbf{z})||$, we also have that,
[…]
We can invoke Lemma B.6 to say
[…]
since $\mathbf{P}^{T}$ is a linear operator, and $\phi(\mathbf{z}^{*})$ satisfies the conditions in the lemma. Therefore, we have,
[…]
We assign terms,
[…]
and once again invoke the Neyman-Fisher Factorization theorem to show $\mathbf{z}^{*}$ is sufficient for $\mathbf{y}$. Since $\epsilon_{\mathbf{y}}$$_{*}$ is a bijective mapping from $\mathbf{z}^{*}$, we have that $\epsilon_{\mathbf{y}}$$_{*}$ is sufficient, and similarly to Theorem 3.2 we state, $p(\mathbf{y}|{\epsilon_{\mathbf{y}}}_{*})=p(\mathbf{y}|{\epsilon_{\mathbf{y}}}_{*},\mathbf{x}_{0})=p(\mathbf{y}|\mathbf{x}_{0})$. ∎
Finally, we note that this proof provides necessary conditions, but not sufficient conditions for the sufficiency of DCS’s estimator. In this work, we do not investigate operators outside of the scope of Theorem B.5, there are potentially even weaker conditions on $\mathcal{A}$ that exist.
Appendix C Invertibility of $\mathcal{A}$
Often, $\mathcal{A}$ is simply non-invertible (e.g. for super-resolution, inpainting, phase retrieval, and sparse MRI reconstruction tasks). With other tasks such as signal deblurring, the invertibility of $\mathcal{A}$ is often mathematically possible, but not numerically stable. In theory, blurring operator can be represented as convolution operators on the signal $\mathbf{x}$. Theorefore, the convolution theorem tells us that inverting a blurring operator $\mathcal{G}(\ast)$ on $\mathbf{x}$ is as simple as taking the quotient of the convolved signal $\mathbf{y}=\mathcal{A}(\mathbf{x})$ against the convolution kernel in the frequency domain, i.e.,
$\mathbf{x}=\mathcal{F}^{-1}[\mathcal{F}(\mathbf{y})/\mathcal{F}(\mathcal{G})]=\mathbf{y}\ast\mathcal{F}^{-1}[\mathcal{F}(\mathcal{G})^{-1}]$ (61)
where $\mathcal{F}$ denotes the Fourier operator. However, in practice, there are implicit assumptions in Eq. 61, such as the computability of $\mathcal{F}(\mathcal{G})$ and the existence of $\mathcal{F}(\mathcal{G})^{-1}$, that may not always hold. In particular, blur kernels are often truncated in practice, resulting in a highly ill-conditioned $\mathcal{F}(\mathcal{G})$ in the frequency domain, and numerical unstable (or non-existent) inverses. Ultimately, directly inverting $\mathcal{A}$ often fails to produce the highest quality results, even though it may be possible.
Appendix D Additional Experiments
In this section, we provide further comparisons against latent and Jacobian-free methods (Table 3).
D.1 Comparison against Latent Models (Table 4)
We show that our pixel-based model also performs favorably against latent models in Table 4. We retain the same experimental setting on pixel-based models as in Table 2. For FFHQ, we use the pretrained FFHQ model weights from (Chung et al. 2022a) for our method, and the pretrained FFHQ model with a VQ-F4 first stage model (Rombach et al. 2022) in latent space models. For ImageNet, we again use pretrained model weights from (Chung et al. 2022a) in pixel-based diffusion solvers, and the Stable Diffusion v1.5 latent model for latent solvers. As with pixel-based methods many existing works suffer in the presence of additional noise. Further implementation details are discussed in Appendix E.
Table 3: Description of latent and Jacobian-free solvers used for comparisons in text. For each solver we list the type (as described in Section A.2), optimization space (pixel or latent), whether it requires backpropagation through a neural function evaluation (NFE, i.e., the score network call), as well as runtime and memory footprint.
Solver Type Space
No NFE
Backprop
Runtime Memory
DCS (Ours) Hybrid Pixel 1x 1x
Latent-DPS (Chung et al. 2023)  [^3] Latent-DPS is a direct application of DPS (Chung et al. 2023) to latent diffusion models. It is also mentioned in (Rout et al. 2023). Posterior Latent $6.1\mathbf{x}$ $8.9\mathbf{x}$
PSLD (Rout et al. 2023) Posterior Latent $7.5\mathbf{x}$ $15\mathbf{x}$
STSL (Rout et al. 2024) Posterior Latent $1.85\mathbf{x}$ $9\mathbf{x}$
ReSample (Song et al. 2024) Projection Latent  [^4] As described in (Song et al. 2024), ReSample does not run backpropagation on the score network, however the implementation does (Appendix E.4). $29.5\mathbf{x}$ $8.95\mathbf{x}$
DPS-JF (Chung et al. 2023) Posterior Pixel $5\mathbf{x} 1.1$$1\mathbf{x}$
LGD-MC (n=10) (Song et al. 2023b) Posterior Pixel 6x 3.2x
LGD-MC-JF (n=10) (Song et al. 2023b) Posterior Pixel $2\mathbf{x}$ $1.1\mathbf{x}$
Table 4: Quantitative comparison against latent models on FFHQ 256x256-1K and ImageNet-1K datasets across various inverse problem tasks and noise levels ($\sigma_{\mathbf{y}}\in\{0.01,0.1\}$).
FFHQ SR $\times 4$ Random Inpainting Box Inpainting Gaussian Deblurring Motion Deblurring $\sigma_{\mathbf{y}}=0.01$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ Ours 0.137 30.138 19.45 0.024 34.839 21.19 0.088 25.112 19.25 0.103 28.688 22.62 0.087 29.480 26.67 Latent-DPS 0.324 20.086 100.27 0.249 22.64 297.43 0.227 22.184 211.23 0.390 25.608 321.5 0.950 -6.753 354.95 PSLD 0.311 20.547 42.26 0.250 22.84 214.08 0.221 22.23 204.87 0.200 23.77 318.20 0.213 23.277 359.40 STSL 0.614 16.063 327.38 0.476 17.859 190.64 0.436 11.843 190.64 0.583 15.196 364.07 0.604 10.095 388.68 ReSample 0.221 24.699 48.87 0.467 22.488 96.89 0.247 20.852 50.3 0.191 27.151 46.5 0.281 25.138 65.06 FFHQ SR $\times 4$ Random Inpainting Box Inpainting Gaussian Deblurring Motion Deblurring $\sigma_{\mathbf{y}}=0.1$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ Ours 0.1748 24.879 30.107 0.1490 27.536 32.800 0.1631 23.217 26.444 0.1763 25.955 26.083 0.2238 24.612 31.400 Latent-DPS 0.3444 19.971 45.052 0.4455 18.117 109.83 0.6410 11.365 326.75 0.6398 13.762 330.93 0.6360 12.524 334.43 PSLD 0.3481 19.251 47.864 0.3105 20.588 41.737 0.3121 19.874 40.428 0.2897 21.068 36.600 0.3307 19.224 40.374 STSL 0.3161 20.279 40.163 0.3722 19.247 54.648 0.5481 13.864 183.00 0.5137 16.411 169.32 0.5188 15.463 163.65 ReSample 0.2613 24.184 50.224 0.5267 21.575 103.62 0.2789 20.581 53.263 0.2984 23.980 56.489 0.6456 19.912 110.42 ImageNet SR $\times 4$ Random Inpainting Box Inpainting Gaussian Deblurring Motion Deblurring $\sigma_{\mathbf{y}}=0.01$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ Ours 0.238 23.452 39.41 0.142 26.063 34.46 0.230 20.625 37.11 0.253 24.218 38.96 0.203 24.619 38.63 Latent-DPS 0.642 17.973 144.82 0.603 19.881 144.81 0.751 11.964 138.33 0.805 10.532 139.62 0.821 10.697 150.49 PSLD 0.380 22.690 168.08 0.306 24.167 125.25 0.330 18.290 156.30 0.397 23.076 134.18 0.453 21.576 187.21 STSL 0.617 19.682 143.62 0.599 20.500 137.09 0.832 9.560 170.93 0.869 8.708 183.38 0.882 8.527 195.74 ReSample 0.552 20.260 133.42 0.820 17.775 229.82 0.504 16.795 138.97 0.513 21.578 116.04 0.573 20.430 145.67 ImageNet SR $\times 4$ Random Inpainting Box Inpainting Gaussian Deblurring Motion Deblurring $\sigma_{\mathbf{y}}=0.1$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ Ours 0.4015 22.988 48.211 0.1655 26.043 34.469 0.2428 19.697 46.026 0.4068 22.283 51.131 0.4348 20.428 61.48 Latent-DPS 0.7257 15.676 147.65 0.7973 9.4153 146.69 0.7980 9.3345 146.51 0.7988 9.3032 193.84 0.8525 9.1369 170.08 PSLD 0.4731 20.875 130.99 0.6068 19.668 145.51 0.7028 13.909 146.74 0.7372 14.181 139.90 0.7504 13.767 149.75 ReSample 0.6514 18.997 155.26 0.9654 13.612 281.82 0.5980 15.843 168.06 0.6814 19.233 173.72 1.0461 15.249 223.52
D.2 Comparison against other Jacobian-Free Methods (Table 5)
A major advantage of DCS is the fact that it is Jacobian-free (Section 3.4) — this results in at least $6\times$ reduction in memory cost during inference compared with Jacobian-based methods, which can be a major enabling factor for the adoption of such algorithms on consumer GPUs and edge devices. However, naively removing the backpropagation through the score network can reduce the quality of the measurement consistency correction step in inverse solvers. In this experiment, we demonstrate that our treatment via the maximum likelihood framework and the noise-aware maximization results in significantly higher quality samples, compared to a naive implementation in DPS-JF and LGD-JF, which are both Jacobian-free variants of the original algorithms (Chung et al. 2022a) and (Song et al. 2023a). Namely, we approximate the Jacobian with respect to the input to the denoising network (left hand side) by the Jacobian with respect to the predicted $\mathbf{x}_{0}$ (right hand side)
[…]
where $\hat{\mathbf{x}}=\mathbf{f}(\mathbf{x}_{t},\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t))$ and $\mathbf{f}$ is an algorithm-dependent function of $\mathbf{x}_{t}$ and its score. (Note that the right hand side no longer involves backpropagation through $\mathbf{f}$ and therefore $\bm{\epsilon}_{\theta}$).
Table 5: Quantitative comparison against other Jacobian-free methods on FFHQ 256x256-1K and ImageNet-1K datasets across various inverse problem tasks and noise levels ($\sigma_{\mathbf{y}}\in\{0.01,0.1\}$).
FFHQ SR $\times 4$ Random Inpainting Box Inpainting Gaussian Deblurring Motion Deblurring $\sigma_{\mathbf{y}}=0.01$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ Ours 0.137 30.138 19.45 0.024 34.839 21.19 0.088 25.112 19.25 0.103 28.688 22.62 0.087 29.480 26.67 DPS-JF 0.488 14.193 44.98 0.335 19.566 58.45 0.178 20.118 28.10 0.211 23.063 34.42 0.289 19.927 40.94 DPS-JF $(T=100)$ 0.589 9.473 41.24 0.578 10.072 42.06 0.571 10.618 43.08 0.563 10.859 43.77 0.566 10.922 41.26 LGD-MC-JF 0.566 10.502 41.25 0.537 12.154 43.85 0.497 13.811 46.40 0.452 15.569 46.22 0.457 15.466 46.08 LGD-MC-JF $(T=100)$ 0.593 9.346 40.60 0.587 9.688 40.99 0.581 10.126 42.30 0.574 10.273 40.59 0.574 10.364 40.51 DDNM 0.208 26.277 51.33 0.040 33.076 23.35 0.209 18.118 88.32 0.235 26.086 71.47 0.424 14.221 250.92 DDRM 0.502 13.002 222.45 0.393 15.935 163.91 0.472 12.148 209.18 - - - - - - FFHQ SR $\times 4$ Random Inpainting Box Inpainting Gaussian Deblurring Motion Deblurring $\sigma_{\mathbf{y}}=0.1$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ Ours 0.1748 24.879 30.107 0.1490 27.536 32.800 0.1631 23.217 26.444 0.1763 25.955 26.083 0.2238 24.612 31.400 DPS-JF 0.494 14.111 46.59 0.371 18.310 56.49 0.226 19.451 34.02 0.246 21.808 35.53 0.342 18.339 40.70 DPS-JF $(T=100)$ 0.589 9.432 40.82 0.582 9.900 39.58 0.572 10.552 42.90 0.564 10.894 42.36 0.568 10.943 42.44 LGD-MC-JF 0.557 11.208 44.86 0.511 13.265 49.07 0.452 15.243 48.68 0.396 17.434 46.76 0.400 17.301 45.53 LGD-MC-JF $(T=100)$ 0.594 9.324 41.06 0.589 9.655 41.65 0.580 10.107 42.97 0.578 10.334 41.84 0.574 10.312 41.53 DDNM 0.6230 21.493 145.889 0.179 24.964 39.183 0.334 19.195 72.105 1.220 10.727 176.756 0.739 5.099 524.021 DDRM 0.7853 6.3273 271.70 0.6018 10.995 255.95 0.6323 9.6360 288.11 - - - - - -
D.3 Further Noise Experiments
Table 6: Quantitative experiments on FFHQ 256x256-1K at $\sigma_{\mathbf{y}}=0.5$. We compare against pixel-based solvers (upper half) and latent-based solvers (lower half).
FFHQ SR $\times 4$ Random Inpainting Box Inpainting Gaussian Deblurring Motion Deblurring LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ Ours 0.2287 20.362 100.94 0.2067 22.999 89.312 0.2005 21.298 40.099 0.2109 24.009 82.132 0.2301 22.306 90.403 DPS 0.2000 22.588 92.791 0.2290 22.808 90.739 0.2118 20.278 81.491 0.2268 25.020 83.686 0.2479 20.767 91.972 DDNM 0.7812 9.8324 387.43 0.8721 15.573 233.15 0.9966 12.607 287.79 1.4475 3.5686 408.85 1.3328 3.1782 393.24 ReSample 0.5704 19.948 179.35 0.6892 20.014 200.06 0.4958 17.530 160.47 0.5409 21.166 162.40 0.6380 19.875 194.69
Figure 9: A demonstration of our solver, DCS, solving two inverse problems on natural images from the CelebA-HQ dataset. Motion blur (left), and box dropout (right) are examples of forward operators that are non-invertible. We show further results in Section 4
D.4 Subset of FFHQ used in other works
Table 7: Quantitative evaluation of our method on FFHQ 256x256, following the experimental setup of (Song et al. 2024). We compare against pixel-based solvers (upper half) and latent-based solvers (lower half).
SR $\times 4$ Random Inpainting Box Inpainting Gaussian Deblurring Motion Deblurring Cost LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ Time $\downarrow$ Mem. $\downarrow$ Ours 0.074 29.51 0.811 0.052 31.13 0.850 0.102 22.07 0.761 0.078 29.92 0.817 0.051 32.32 0.833 $\mathbf{1x}$ $\mathbf{1x}$ DPS 0.132 27.10 0.729 0.084 30.91 0.833 0.107 21.62 0.755 0.090 28.26 0.767 0.108 26.816 0.726 $6\mathbf{x}$ $3.2\mathbf{x}$ MCG 0.112 27.07 0.784 0.877 11.02 0.02 0.905 10.883 0.001 0.176 24.89 0.768 - - - $6.1\mathbf{x}$ $3.2\mathbf{x}$ DDNM 0.242 27.63 0.587 0.230 27.92 0.604 0.194 23.08 0.639 0.287 27.24 0.561 0.642 8.682 0.165 $1.75\mathbf{x}$ $1\mathbf{x}$ Latent-DPS 0.324 20.086 0.473 0.249 22.64 0.570 0.227 22.184 0.595 0.209 23.512 0.600 0.217 22.930 0.582 $6.1\mathbf{x}$ $8.9\mathbf{x}$ PSLD 0.311 20.547 0.491 0.250 22.84 0.579 0.221 22.23 0.607 0.200 23.77 0.614 0.213 23.277 0.596 $7.5\mathbf{x}$ $15\mathbf{x}$ STSL 0.242 27.63 0.587 0.230 27.92 0.604 0.194 23.08 0.639 0.287 27.24 0.561 0.641 10.17 0.245 $1.85\mathbf{x}$ $9\mathbf{x}$ ReSample 0.090 29.024 0.791 0.053 30.99 0.844 0.156 20.71 0.778 0.113 29.19 0.784 0.197 27.65 0.706 $29.5\mathbf{x}$ $8.95\mathbf{x}$
Appendix E Implementation Details
We provide implementation details of our experiments, as well as those for other experiments we compare against.
E.1 Our Method
Our proposed DCS has just two primary hyperparameters, as described in the table below. First is the number of time steps $T$. This has relatively little effect on our model performance on most tasks. However, it is occasionally helpful to increase $T$, especially in box inpainting, where there is zero signal from $\mathbf{y}$ in the masked region. Here, higher $T$ allows the diffusion model to obtain a better solution in this unconditional diffusion process. Second, we have the choice of minimizer, which is by default the Adam optimizer (Kingma and Ba 2014). However, in the case of linear $\mathcal{A}$, this optimizer can be replaced by the closed form analytical solution to $\mathcal{A}(\mathbf{x})=\mathbf{y}$.
For nearly all experiments, we use the Adam optimizer with 50 optimization steps and a learning rate of $1$. The exceptions are the random inpainting and box inpainting tasks, where there is no conditioning information on the masked pixels. This requires more denoising steps, as the diffusion process is totally unconditional inside the mask, up to local correlations learned inside the score network $s_{\theta}$. Here, we use the analytical solver with $\mathcal{A}^{\dagger}=\mathcal{A}$. Similarly, for nearly all experiments we use $T=50$ as found in Table 8, with the exception being random inpainting and box inpainting tasks, where we found that taking $T=1000$ steps improved performance. However, there is little increase in runtime, since the minimization step is much faster here.
Notation            Definition
$T$ The number of diffusion steps used in the sampler.
minimizer The minimizer used to solve for $\epsilon_{\mathbf{y}}$}.
E.2 Latent Models on ImageNet
We note that previous latent models use the pretrained weights in (Rombach et al. 2022) for $256\times 256$ resolution datasets. However, there are no published weights in the GitHub repository for unconditional ImageNet, making a fair comparison of our method against latent models more involved. To this end, we leverage a significantly more powerful Stable Diffusion v1.5 model, with publicly available weights on HuggingFace for our experiments. The measurements and the output images are appropriately scaled for a fair comparison.
E.3 STSL
At the time of writing this work, we did not find publicly available code for STSL (Rout et al. 2024). Therefore, we implement the algorithm ourselves in our codebase, and use the hyperparameters provided in the paper.
E.4 ReSample
We directly use the published code of ReSample (Song et al. 2024) with no changes in our paper. We discuss two notable aspects of the experiments with ReSample. First, the implementation on GitHub differs from that pseudocode discussed in the paper. Namely, the pseudocode in the paper describes enforcing latent- and pixel-based consistency occasionally during an otherwise unconditional sampling process.
In the code we observed that the sampling step taken is actually a DPS (Chung et al. 2022a) sampling step, which includes a posterior-based guidance step that takes an expensive gradient of the noise function. To see this, note that L255 in the resample_sampling function in ddim.py calls a function measurement_cond_fn, which is defined at L62 in main.py and passed into the resampling function. This function is a member of the class PosteriorSampling defined in L53 in condition_methods.py. Inspecting this class, we note that it calls torch.autograd.grad on the diffusion step as a function of x_prev (L33 or L39). In other words, a gradient is computed for the measurement norm with respect to the input to the diffusion model, i.e., a DPS step.
We closely investigated this DPS step in our experiments, ultimately concluding that it has a significant effect on the performance of the algorithm, and that it was a more fair comparison to include this step, rather than removing it. However, the inclusion of this sampling step has two primary effects. First, it results in further increases the computation time of ReSample. Second it reveals that ReSample relies significantly on a posterior-based formulation, applying additional resampling steps at each stage.
In experiments, we note that ReSample is significantly slower than other algorithms during sampling (see Table 1). For example, sampling $\sim 1000$ images with ImageNet takes more than two weeks on an A6000 GPU. Since we run five different experimental conditions for each dataset, this was an unacceptably long runtime for our academic resources. Therefore, we reduce the number of diffusion steps $T$ of ReSample in our experiments, from $500$ reported in (Song et al. 2024) to $50$. However, we do provide a single experiment from the (Song et al. 2024) paper, where we reproduce the hyperparameters and dataset (a $100$ image subset of FFHQ). We note that (Song et al. 2024) took a subset of the FFHQ dataset, where performance differed from the full $\times256$-1K dataset performance (c.f. Table 2). Since the subset was not published, we selected a dataset based where ReSample obtained the same performance with its default parameters in (Song et al. 2024) (Table 7).
E.5 DDRM
We used the version of DDRM which is implemented in the DDNM codebase. While DDRM may theoretically be able to handle deblurring tasks, due to the high rank of the forward operators, the SVD cannot be explicitly defined in memory, and no existing code-base for DDRM supplies fast and memory-saving versions of these operators. Because in our settings DDRM’s performance was similar or worse than that of DDNM, and due to the fact that DDRM can be considered a subtype of DDNM (see Appendix of (Wang, Yu, and Zhang 2022)), we do not run benchmarks.
Appendix F Further Qualitative Comparisons
We provide further qualitative examples from the FFHQ $\times256$-1K and ImageNet $\times256$-1K datasets accompanying our quantitative evaluation in Table 2.
[…]
Figure 11: Comparison against competing works on FFHQ $\times256$-1K dataset with the random inpainting task.
Figure 12: Comparison against competing works on FFHQ $\times256$-1K dataset with the box inpainting task.
Figure 13: Comparison against competing works on FFHQ $\times256$-1K dataset with the Gaussian deblurring task.
Figure 14: Comparison against competing works on FFHQ $\times256$-1K dataset with the motion deblurring task.
[…]
Figure 16: Comparison against competing works on ImageNet $\times256$-1K dataset with the random inpainting task.
Figure 17: Comparison against competing works on FFHQ $\times256$-1K dataset with the box inpainting task.
Figure 18: Comparison against competing works on ImageNet $\times256$-1K dataset with the Gaussian deblurring task.
Figure 19: Comparison against competing works on ImageNet $\times256$-1K dataset with the motion deblurring task.
Figure 20: Comparison against competing works on FFHQ $\times256$-1K dataset with the random inpainting task at various noise levels.
Figure 21: Comparison against competing works on FFHQ $\times256$-1K dataset with the box inpainting task at various noise levels.
Figure 22: Comparison against competing works on FFHQ $\times256$-1K dataset with the Gaussian deblurring task at various noise levels.
Figure 23: Comparison against competing works on FFHQ $\times256$-1K dataset with the motion deblurring task at various noise levels.